# Large Language Models can be Guided to Evade AI-Generated Text Detection

### v5

@@ Abstract
Large language models (LLMs) have shown remarkable performance in various tasks and have been extensively utilized by the public. However, the increasing concerns regarding the misuse of LLMs, such as plagiarism and spamming, have led to the development of multiple detectors, including fine-tuned classifiers and statistical methods. In this study, we equip LLMs with prompts, rather than relying on an external paraphraser, to evaluate the vulnerability of these detectors. We propose a novel Substitution-based In-Context example Optimization method (SICO) to automatically construct prompts for evading the detectors. SICO is cost-efficient as it requires only 40 human-written examples and a limited number of LLM inferences to generate a prompt. Moreover, once a task-specific prompt has been constructed, it can be universally used against a wide range of detectors. Extensive experiments across three real-world tasks demonstrate that SICO significantly outperforms the paraphraser baselines and enables GPT-3.5 to successfully evade six detectors, decreasing their AUC by 0.5 on average. Furthermore, a comprehensive human evaluation as well as a validation experiment in the wild show that the SICO-generated text achieves human-level readability and task completion rates. Finally, the strong performance of SICO exhibits its potential as a reliable evaluation tool for future detectors. The codes and data are located on https://github.com/ColinLu50/Evade-GPT-Detector.
@@ 1 Introduction
The rapid advancement of large language models (LLMs), such as GPT (Brown et al., 2020), PaLM (Chowdhery et al., 2022), and LLaMa (Touvron et al., 2023), has led to a largely-increased capacity for generating high-quality human-like text. However, there are also growing concerns surrounding the misuse of these models, including generating fake product reviews (Adelani et al., 2020; Lin et al., 2022) and misinformation (Lin et al., 2022), enabling academic dishonesty (Stokel-Walker, 2022), and producing misleading answers on websites (StackOverflow, 2023).
@@
We assess the evasion performance of SICO across three real-world tasks that are susceptible to the misuse of LLMs, i.e., academic essay writing, open-ended question answering, and fake review generation. The results demonstrate that SICO consistently outperforms the paraphraser baselines, leading to a decrease in AUC by approximately 0.5 on average for six existing detectors. Additionally, a comprehensive human evaluation involving 600 examples shows that the SICO-generated text is comparable to, and in some cases even better than, human-written text in terms of readability and task completion rates. To further evaluate the practical utility of SICO, we deploy it on Reddit, an online social platform, to generate responses for users’ questions. The high percentage of generated responses that are liked by Reddit users shows that SICO is capable of generating human-approved content while being barely identified as AI. In addition to its strong evasion performance, SICO is also cost-efficient and easy to use. Unlike paraphraser-based methods that often require extensive computational resources – as evidenced by the fine-tuning of a 13B model on a large dataset (Krishna et al., 2023) – SICO only requires 40 human-generated examples and a limited number of LLM inferences (e.g., costing approximately 1 USD using the GPT-3.5 API). Besides, once a task-specific prompt has been constructed by SICO, it can be universally used against a wide range of detectors.
Considering the importance of detecting AI-generated text to avoid their misuse, the results presented in this work certainly reveal the vulnerability of the existing detectors. Besides, this work presents the first empirical evidence that LLMs can evade detectors through a prompt-guided approach. Finally, the strong evasion performance of SICO suggests that it can be used as a standard evaluation tool for any future AI-generated text detectors. We hope that these findings can better facilitate the research concerning the responsible use of LLMs. To summarize, our main contributions are:
@@
  • A comprehensive human evaluation, as well as a validation experiment in the wild, verifies that the SICO-generated text achieves human-level readability and task completion rates.
@@ 2.1 AI-generated text detection
In recent years, the research community has developed a wide range of detectors for AI-generated contents. In general, these detectors can be classified into three categories: training-based, statistical, and watermarking methods. Training-based methods treat the detection problem as a binary classification task, where neural networks are trained using AI-generated text and human-written text. Early studies utilized classifiers to identify fake reviews (Hovy, 2016) and fake news (Zellers et al., 2019). More recently, researchers have trained classifiers using text generated by LLMs, such as the GPT-3.5 detector (Guo et al., 2023) and GPT-2 detector (Solaiman et al., 2019).
Statistical methods, on the other hand, focus on zero-shot detection without any additional training overhead. These methods seek to distinguish between human-written text and AI-generated text based on the statistical characteristics of text, such as the statistical irregularities in measures like entropy (Lavergne et al., 2008), perplexity (Beresneva, 2016) and token rank (Gehrmann et al., 2019). A recent method, DetectGPT (Mitchell et al., 2023), exploits the phenomenon that AI-generated text tends to lie in the negative curvature regions of log probability of text.
The watermarking methods involve modifying the LLM’s text generation process to imprint specific patterns on the generated text, such that it can be detected (Abdelnabi and Fritz, 2021; Grinbaum and Adomaitis, 2022; Kirchenbauer et al., 2023). Although the proposed method SICO primarily focuses on the first two types of detection methods, it can also help evade watermarking when acted as an external paraphraser, as shown in Appendix G.
@@ 2.2 In-context learning
With the increasing scales of models and corpora (Devlin et al., 2019; Radford et al., 2019; Chowdhery et al., 2022; Gou et al., 2022), LLMs have demonstrated the in-context learning (ICL) ability, allowing them to perform tasks with only a few examples provided as demonstrations (Brown et al., 2020). Recent studies have focused on designing demonstrations during inference, which can be divided into demonstration selection, ordering, and formatting (Dong et al., 2022). Specifically, demonstrations can be selected based on unsupervised metrics or supervised strategies (Kim et al., 2022; Gonen et al., 2022; Rubin et al., 2022). For ordering, Liu et al. (2021) sorted examples by their distances to the input. Regarding demonstration formatting, Wei et al. (2022) proposed the so-called chain-of-thoughts (COT) format, and subsequent works have developed automatic COT (Zhang et al., 2022). In contrast to these works, we focus on iteratively optimizing demonstrations through substitutions. In principle, the proposed method SICO can be used in combination with the above-mentioned methods, potentially leading to improved performance.
@@ 3.2 Prompt Construction
Feature extraction  This step involves the $K$ pairs of AI-generated and human-written outputs from $D$, denoted by $\{(y_{\text{AI}}^{k},y_{\text{human}}^{k})\}_{k=1}^{K}$. We provide LLM with these pairs and ask LLM to extract the distinct linguistic features of human-written text, denoted as $t_{\text{feature}}$.
@@
In-context example optimization  We initialize the in-context examples as $(x_{\text{ic}}^{k},y_{\text{ic}}^{k})$, where $y_{\text{ic}}^{k}$ is generated by paraphrasing $y_{\text{AI}}^{k}$. More specifically, the text feature $t_{\text{feature}}$ is concatenated with a paraphrasing instruction to instruct LLM to paraphrase $y_{\text{AI}}^{k}$ to generate $y_{\text{ic}}^{k}$
Then the in-context output $y_{\text{ic}}$ is iteratively optimized to be less AI-like, which is directed by the proxy detector. By presenting more and more representative in-context demonstrations to LLM, it is expected to understand how to generate human-like outputs. Formally, the optimization goal can be expressed as:
where $\text{SIM}(y_{\text{ic}})$ denotes the set of text that is semantically similar to $y_{\text{ic}}$. The goal of setting such semantic restriction is to maintain the usability of the text during optimization. In SICO, we generate semantically similar text by replacing words with their synonyms and rephrasing sentences. This is explained in detail below.
Substitution type
To generate $y^{\prime}_{\text{ic}}$ that is semantically similar to $y_{\text{ic}}$, we employ substitution at word level and sentence level in turn. For word-level substitution, we use WordNet (Miller, 1998), a lexical database of English words, to construct a synonym substitution set. We restrict substitutions to content words that carry meanings and ensure that the substitution would not change the part-of-speech tags. We use a mask language model to filter out the candidate words that not fits the context. For sentence-level substitution, we utilize a paraphrasing instruction combined with extracted feature, denoted as $t_{\text{feature}}\oplus p_{\text{para}}$. This combined instruction is used to prompt LLM to generate paraphrases for each sentence in $y_{\text{ic}}$.
Algorithm  As illustrated in Algorithm 1, SICO would optimize $\{y_{\text{ic}}^{k}\}_{k=1}^{K}$ for $N$ iterations (lines 4-17). At each iteration, each $y_{\text{ic}}^{k}$ would be optimized by greedy substitution (line 11), as presented in Algorithm 2. Specifically, for the $i$-th original word/sentence $y_{i}$ in the text $y$, let $C_{i,j}$ denote its $j$-th synonym/paraphrase, and let $\text{SUB}(y_{i},C_{i,j})$ denote the new text resulting from substituting $y_{i}$ with $C_{i,j}$. For each $y_{i}$, SICO finds the best synonym/paraphrase $C_{i,*}$ by checking which $C_{i,j}$ gives the lowest AI-probability when substituting $y_{i}$ (Line 1 in Algorithm 2).
@@ 4.1 Experimental setup
For academic writing, we employ Wikipedia paragraphs from SQuAD dataset (Rajpurkar et al., 2016) as human-written text. Following the approach in Mitchell et al. (2023), we use the first 30 words of these paragraphs as task inputs and ask GPT-3.5 to complete the rest. For open-ended question answering, we sample questions from Eli5 (Fan et al., 2019) dataset and ask GPT-3.5 to generate answers, following Krishna et al. (2023). For fake review generation, we first instruct GPT-3.5 to extract the business name and five keywords from human-written reviews from Yelp dataset (Zhang et al., 2015), and then generate fake reviews based on the extracted information with specified sentiment. For each task, we collect 200 examples from GPT-3.5 (called original AI-generated text) and 200 human-written examples from corresponding dataset. More details about dataset can be found in Appendix F.
@@
Figure 2: ROC curves of six detectors on the text generated by different evasion methods on academic writing task.
@@ 4.2 Evasion Performance and Analysis
Table 1 presents the performance of SICO and other baselines against six detectors in AUC score. SICO consistently outperforms other baselines by a substantial margin in all cases. Notably, in most cases, SICO reduces the AUC score to less than 0.5, equivalent to the expected performance of a random classifier. Figure 2 shows the ROC curves of evasion methods on academic writing task. One can clearly observe that SICO curves lie below others along different thresholds, often lower than the random classifier curve. More evasion results including ROC cures and detection rates are shown in Appendix H.
@@ 4.3 Human Evaluation
4.4 Real-life Experiments
To further assess the applicability of SICO in real-world, we simulate one potential misuse case of LLM, where SICO is deployed as an automatic reply bot on Reddit [^2] https://www.reddit.com/, a popular online social platform. We wrote a script to monitor the new posts submitted in the community for asking questions and used GPT-3.5 equipped with SICO-Para to automatically reply them. The prompt we used is trained for question answering task. On Reddit, except for giving comments, users can express their approval of other’s responses by clicking the “like” or “dislike” button. To minimize the social impact, we limit the number of responses to 40 and deleted them after collecting results.
The quantitative results in Table 3 demonstrate that users generally react positively to the text from SICO. Specifically, 40% of the responses from SICO receive “likes” from Reddit users, significantly higher than the 2.5% that are disliked. The remaining 57.5% of responses go unnoticed, which is common in social media. Besides, in 12.5% of cases, users left additional comments to express their approval or to ask further questions, demonstrating their engagement and interest in SICO’s response. Figure 3 presents two SICO’s responses that got approval from users, as indicated by “likes” and comments.
Table 3: Quantitative results of Reddit test.
Like % Dislike % Comment %
40% 2.50% 12.5%
@@ 4.5 Cost Efficiency
Figure 3: Screenshots of replies from SICO that got approved by Reddit users. (Left) Approval is shown by the number of “like”s. (Right) The original poster express thanks to SICO’s response.
@@ 5.2 SICO with Different Proxy Detectors and LLMs
As described in Section 3, SICO requires a proxy detector and a LLM to construct a prompt. In this experiment, we explore the performance of SICO-Para on writing task, using three types of proxy detectors: (1) training-based model GPT-3.5 detector, (2) API detector GPTzero, and (3) statistical method DetectGPT. For different LLMs, we adopt Vicuna-13B (Chiang et al., 2023), an open-source chatbot fine-tuned from LLaMa (Touvron et al., 2023). Results in Table 5 show that SICO maintains a high degree of detection evasion performance, regardless of proxy detectors or LLMs. In most cases, SICO manages to reduce the AUC of detectors by approximately 0.4.
@@
5.3 Examples of the generated text
Table 6 presents the fake reviews generated by SICO-Gen and SICO-Para. The generated text shows high readability and fulfill the task’s goal, successfully mentioning all keywords and generating positive reviews of the specified object. The AI probability, denoted as $\mathcal{P}_{\text{AI}}$ in the table, is determined by the GPT3-D. More examples are shown in Appendix I.

### v6

@@ Abstract
Large language models (LLMs) have shown remarkable performance in various tasks and have been extensively utilized by the public. However, the increasing concerns regarding the misuse of LLMs, such as plagiarism and spamming, have led to the development of multiple detectors, including fine-tuned classifiers and statistical methods. In this study, we equip LLMs with prompts, rather than relying on an external paraphraser, to evaluate the vulnerability of these detectors. We propose a novel Substitution-based In-Context example Optimization method (SICO) to automatically construct prompts for evading the detectors. SICO is cost-efficient as it requires only 40 human-written examples and a limited number of LLM inferences to generate a prompt. Moreover, once a task-specific prompt has been constructed, it can be universally used against a wide range of detectors. Extensive experiments across three real-world tasks demonstrate that SICO significantly outperforms the paraphraser baselines and enables GPT-3.5 to successfully evade six detectors, decreasing their AUC by 0.5 on average. Furthermore, a comprehensive human evaluation show that the SICO-generated text achieves human-level readability and task completion rates, while preserving high imperceptibility. Finally, we propose an ensemble approach to enhance the robustness of detectors against SICO attack.  [^1] The code is publicly available at https://github.com/ColinLu50/Evade-GPT-Detector.
@@ 1 Introduction
The rapid advancement of large language models (LLMs), such as GPT (Brown et al., 2020) and LLaMa (Touvron et al., 2023), has led to a largely-increased capacity for generating high-quality human-like text. However, there are also growing concerns surrounding the misuse of these models, including generating fake product reviews (Adelani et al., 2020; Lin et al., 2022) and misinformation (Lin et al., 2022), enabling academic dishonesty (Stokel-Walker, 2022), and producing misleading answers on websites (StackOverflow, 2023).
@@
We assess the evasion performance of SICO across three real-world tasks that are susceptible to the misuse of LLMs, i.e., academic essay writing, open-ended question answering, and fake review generation. The results demonstrate that SICO consistently outperforms the paraphraser baselines, leading to a decrease in AUC by approximately 0.5 on average for six existing detectors. Additionally, a comprehensive human evaluation involving 600 examples shows that the SICO-generated text is comparable to, and in some cases even better than, human-written text in terms of readability and task completion rates. It also demonstrates that SICO reduces the probability of being recognized by humans. In addition to its strong evasion performance, SICO is also cost-efficient and easy to use. Unlike paraphraser-based methods that often require extensive computational resources – as evidenced by the fine-tuning of a 13B model on a large dataset (Krishna et al., 2023) – SICO only requires 40 human-generated examples and a limited number of LLM inferences (e.g., costing approximately 1 USD using the GPT-3.5 API). Besides, once a task-specific prompt has been constructed by SICO, it can be universally used against a wide range of detectors.
Considering the importance of detecting AI-generated text to avoid their misuse, the results presented in this work certainly reveal the vulnerability of the existing detectors. Besides, this work presents the first empirical evidence that LLMs can evade detectors through a prompt-guided approach. The strong evasion performance of SICO suggests that it can be used as a standard evaluation tool for any future AI-generated text detectors. Finally, we propose an ensemble approach to enhance the robustness of detectors against SICO attack. We hope that these findings can better facilitate the research concerning the responsible use of LLMs. To summarize, our main contributions are:
@@
  • A comprehensive human evaluation verifies that the SICO-generated text achieves human-level readability and task completion rates, while preserving high imperceptibility.
@@ 2.1 AI-generated text detection
In recent years, the research community has developed a wide range of detectors for AI-generated contents. In general, these detectors can be classified into three categories: training-based, statistical, and watermarking methods. Training-based methods treat the detection problem as a binary classification task, where neural networks are trained using AI-generated text and human-written text. Early studies utilized classifiers to identify fake reviews (Hovy, 2016) and fake news (Zellers et al., 2019). More recently, researchers have trained classifiers using text generated by LLMs, such as the GPT-3.5 detector (Guo et al., 2023) and GPT-2 detector (Solaiman et al., 2019). Statistical methods, on the other hand, focus on zero-shot detection without any additional training overhead. These methods seek to distinguish between human-written text and AI-generated text based on the statistical characteristics of text, such as the statistical irregularities in measures like entropy (Lavergne et al., 2008), perplexity (Beresneva, 2016) and token rank (Gehrmann et al., 2019). A recent method, DetectGPT (Mitchell et al., 2023), exploits the phenomenon that AI-generated text tends to lie in the negative curvature regions of log probability of text. The watermarking methods involve modifying the LLM’s text generation process to imprint specific patterns on the generated text, such that it can be detected (Abdelnabi & Fritz, 2021; Grinbaum & Adomaitis, 2022; Kirchenbauer et al., 2023). Although the proposed method SICO primarily focuses on the first two types of detection methods, it can also help evade watermarking when acted as an external paraphraser, as shown in Appendix F.
Recent studies have found that paraphrasing can evade these detectors, which trains an additional neural network to rewrite the original AI-generated text (Krishna et al., 2023; Sadasivan et al., 2023). In contrast, SICO eliminates the need for extra models or training steps. SICO provides an automatic approach that iteratively improves prompt, unlike the in-the-wild prompts, which which are typically discovered through manual trial and error (Uploader, 2023).
@@ 2.2 In-context learning
With the increasing scales of models and corpora (Radford et al., 2019; Chowdhery et al., 2022; Gou et al., 2022; Wei et al., 2024), LLMs have demonstrated the in-context learning (ICL) ability, allowing them to perform tasks with only a few examples provided as demonstrations (Brown et al., 2020). Recent studies have focused on designing demonstrations during inference, which can be divided into demonstration selection, ordering, and formatting (Dong et al., 2022). Specifically, demonstrations can be selected based on unsupervised metrics or supervised strategies (Kim et al., 2022; Gonen et al., 2022; Wei et al., 2023). For ordering, Liu et al. (2021) sorted examples by their distances to the input. Regarding demonstration formatting, Wei et al. (2022) proposed the so-called chain-of-thoughts (COT) format, and subsequent works have developed automatic COT (Zhang et al., 2022). In contrast to these works, we focus on iteratively optimizing demonstrations through substitutions. In principle, the proposed method SICO can be used in combination with the above-mentioned methods, potentially leading to improved performance.
@@ 3.2 Prompt Construction
Feature extraction  This step involves the $K$ pairs of AI-generated and human-written outputs from $D$, denoted by $\{(y_{\text{AI}}^{k},y_{\text{human}}^{k})\}_{k=1}^{K}$. We provide LLM with these pairs and ask LLM to extract the distinct linguistic features of human-written text, denoted as $t_{\text{feature}}$. The top left of Figure 1 demonstrates this process and generate text to describe the feature of human-written text. This feature text is then utilized for sentence paraphrasing and included in the final prompt.
In-context example optimization  We initialize the in-context examples as $(x_{\text{ic}}^{k},y_{\text{ic}}^{k})$, where $y_{\text{ic}}^{k}$ is generated by paraphrasing $y_{\text{AI}}^{k}$. More specifically, the feature $t_{\text{feature}}$ is concatenated with a paraphrasing instruction to instruct LLM to paraphrase the AI-generated text to obtain the initial $y_{\text{ic}}^{k}$. The paraphrasing instruction is “Based on the description, paraphrase the following text to be human style”.
Then the in-context output $y_{\text{ic}}$ is iteratively optimized to be less AI-like, as determined by the probability $\mathcal{P}_{\text{AI}}$ calculated by the proxy detector. By presenting more and more representative, i.e. lower $\mathcal{P}_{\text{AI}}$ , in-context demonstrations to LLM, it is expected to understand how to generate human-like outputs. This in-context example optimization procedure is the key step in SICO for improving evasion performance, as verified by the ablation study in Section 5.1. Formally, the optimization goal can be expressed as:
where $\text{SIM}(y_{\text{ic}})$ denotes the set of text that is semantically similar to $y_{\text{ic}}$. The goal of setting such semantic restriction is to maintain the usability of the text during optimization. In SICO, we generate semantically similar text by replacing words and rephrasing sentences. This is explained in detail below.
@@
Substitution type
To generate $y^{\prime}_{\text{ic}}$ that is semantically similar to $y_{\text{ic}}$, we employ substitution at word level and sentence level in turn. For word-level substitution, we use WordNet (Miller, 1998), a lexical database of English words, to construct a synonym substitution set. We restrict substitutions to content words [^2] Content words are the words that carry meanings, consisting of nouns, verbs, adjectives and adverbs. and ensure that the substitution would not change the part-of-speech tags. We use a mask language model to filter out the candidate words that not fits the context. For sentence-level substitution, we utilize the paraphrasing instruction combined with extracted feature, denoted as $t_{\text{feature}}\oplus p_{\text{para}}$. This combined instruction is used to prompt LLM to generate paraphrases for each $y_{\text{ic}}$. Using $t_{\text{feature}}$ will make the generated paraphrases be more human-like, thus increasing the efficiency of optimization.
@@
Algorithm  As illustrated in Algorithm 1, SICO would optimize $\{y_{\text{ic}}^{k}\}_{k=1}^{K}$ for $N$ iterations (lines 4-17). At each iteration, each $y_{\text{ic}}^{k}$ would be optimized by greedy substitution (line 11), as presented in Algorithm 2. Specifically, for the $i$-th original word/sentence $y_{i}$ in the text $y$, let $C_{i,j}$ denote its $j$-th synonym/paraphrase, and let $\text{SUB}(y_{i},C_{i,j})$ denote the new text resulting from substituting $y_{i}$ with $C_{i,j}$. For each substitution position $i$, SICO identifies the best substitution $C_{i,*}$ by checking which $C_{i,j}$ results in the lowest AI probability when replacing $y_{i}$ (Line 1 in Algorithm 2).
@@ 4.1 Experimental Setup
For academic writing, we employ Wikipedia paragraphs from SQuAD dataset (Rajpurkar et al., 2016) as human-written text. Following the approach in Mitchell et al. (2023), we use the first 30 words of these paragraphs as task inputs and ask GPT-3.5 to complete the rest. For open-ended question answering, we sample questions from Eli5 (Fan et al., 2019) dataset and ask GPT-3.5 to generate answers, following Krishna et al. (2023). For fake review generation, we first instruct GPT-3.5 to extract the business name and five keywords from human-written reviews from Yelp dataset (Zhang et al., 2015), and then generate fake reviews based on the extracted information with specified sentiment. For each task, we collect 200 examples from GPT-3.5 (called original AI-generated text) and 200 human-written examples from corresponding dataset. More details about dataset can be found in Appendix E.
@@
Figure 2: ROC curves for six detectors evaluating text generated by various evasion methods in an academic writing task.
@@ 4.2 Evasion Performance and Analysis
Table 1 presents the performance of SICO and other baselines against six detectors in AUC score. SICO consistently outperforms other baselines by a substantial margin in all cases. Notably, in most cases, SICO reduces the AUC score to less than 0.5, equivalent to the expected performance of a random classifier. Figure 2 shows the ROC curves of evasion methods on academic writing task. One can clearly observe that SICO curves lie below others along different thresholds, often lower than the random classifier curve. More evasion results including ROC cures and detection rates are shown in Appendix G.
@@ 4.3.1 Text Quality
4.4 Imperceptibility
Table 3: Imperceptibility of text from different sources.
Source Writing QA Review
AI 24% 52% 54%
DIPPER 31% 46% 66%
SICO 22% 14% 37%
Human 18% 13% 22%
Another key factor for the usability of AI-generated text is its imperceptibility. If text is easily identified as AI-generated by humans, its usability might be significantly influenced. Thus, we conducted an experiment to assess the imperceptibility of text. We sampled 200 examples (50 each from AI, DIPPER, SICO-Gen, and human) across three tasks. Three human annotators were asked to identify whether the text was AI-generated or human-written. All annotators in the experiment used ChatGPT before. The values in Table 3 show the percentage of texts recognized as AI-generated by the annotators, where a lower percentage indicates higher imperceptibility. The results demonstrate that SICO remarkably reduces the probability of being recognized by annotators for QA and review generation tasks. For the academic writing task, the similar low detection rates for AI and human-generated text (24% VS 18%) annotators’ detection capabilities are less accurate for this task, explaining why SICO is less effective here. More details can be found in Appendix D.
@@ 5.2 SICO with Different Proxy Detectors and LLMs
As described in Section 3, SICO requires a proxy detector and a LLM to construct a prompt. In this experiment, we explore the performance of SICO-Para on writing task, using three types of proxy detectors: (1) training-based model GPT-3.5 detector, (2) API detector GPTzero, and (3) statistical method DetectGPT. For different LLMs, we adopt Vicuna-13B (Chiang et al., 2023), an open-source chatbot fine-tuned from LLaMa (Touvron et al., 2023). Results in Table 5 show that SICO maintains a high degree of detection evasion performance, regardless of proxy detectors or LLMs. In most cases, SICO manages to reduce the AUC of detectors by approximately 0.4. More results of using different LLMs can be found in Appendix B.4.
@@
5.3 Examples of SICO text